In a consumer satisfaction survey the responses are “unsatisfactory,” “average,” and “satisfactory”. If the response probability for “unsatisfactory” is 0.21, and the response probability for “average” is 0.41, what is the probability that in surveys the answer for “unsatisfactory” will be selected 3 times, the answer for “average” will be selected 5 times, and the answer for “satisfactory” will be selected 2 times?
Answer #1

here probability for satisfactory =1-0.21-0.41=0.38

therefore P( answer for “unsatisfactory” will be selected 3 times, the answer for “average” will be selected 5 times, and the answer for “satisfactory” will be selected 2 times )

=(10!/(3!*5!*2!))*(0.21)3(0.41)5*(0.38)2 =0.0390